OpenStudy (anonymous):

Find the critical numbers of the function. (Round your answers to three decimal places.) s(t)=3t^4+ 4t^3 − 6t^2 will fan and medal!!!

OpenStudy (anonymous):

\(s'(t)=4.3t^3 + 4.3t^2-6.2t\) Solve \(12t^3+12t^2-12t=0\), you will get \(x=0\), \(x=\frac{-1+\sqrt5}{2} \approx 0.618\), \(x=\frac{-1-\sqrt5}{2} \approx -1.618\)